# Guiding Cross-Modal Representations with MLLM Priors via Preference Alignment

### v1

@@ 4.3 Ablation Studies
OpenCLIP (ViT-H/14[27] 23.64 0.50 47.28 18.05 0.50 36.1
Qwen2-VL-7B [6] 6.89 6.34 1.09 7.63 2.13 3.58
@@ Appendix A Preliminaries
Recent works [31, 32] prompt LLMs to make binary relevance judgments and construct candidate rankings. Inspired by these works, we leverage MLLMs to generate fine-grained alignment scores. For each image-text pair, we construct a specific relevance prompt “<image> Does the image align with the text <text>? Answer Yes or No”. This prompt, containing both the image and the text, is fed into the MLLM. The MLLM then processes the input and generates token logits, including those for the target tokens “Yes” and “No”.
@@ B.1 Detailed Training Data Curation
Our primary training data originates from the large-scale OpenImage dataset [26]. To curate a high-quality dataset suitable for fine-grained learning, we employed the following procedure:
@@
    Clustering and Deduplication: We first applied clustering and deduplication to the Human-verified OpenImage data to reduce redundancy and group similar images. Following the SemDeDup method [23], we set the number of clusters to 50,000 and used an epsilon value of 0.07 to filter out near-duplicate images.
@@
    Hard Negative Mining: For each anchor image in the deduplicated set, we identified hard negative images by retrieving the top-3 most similar images based on their DINOv2 [22] embeddings.
@@
    Caption Generation: For each anchor image and its selected top-3 hard negative images, we employed Qwen2.5-VL-72B [7] to generate corresponding captions for paired images (constructing 6 distinct image pairs from each anchor image and its top-3 hard negative images). To enhance descriptive diversity, we configured the generation temperature to 0.7 and applied the generation process three times, resulting in rich and varied textual representations. As demonstrated in Figure 4, this method enabled us to generate detailed comparative descriptions that effectively capture the visual distinctions between images.
@@ B.2 Detailed Implementation Settings
MAPLE consists of two models: a reward model (Qwen2-VL-7B) and a policy model (initialized with Qwen2-VL-2B or Qwen2-VL-7B). We fine-tune the policy model using Low-Rank Adaptation (LoRA) [33], applying it only to the LLM component’s attention and projection layers, while keeping the vision encoder and connector frozen. We employ a differential learning rate strategy, with the base learning rate applied to the LoRA parameters and amplifying the base learning rate by 100 times for the newly added parameters $\tau$ and $\beta$. Key model-related hyperparameters are summarized in Table 5.
@@
Training was conducted on a cluster of 32 NVIDIA A100 GPUs (80GB memory each). To optimize computational efficiency and memory usage, we employed bfloat16 mixed-precision training, gradient checkpointing for the LLM component, and Flash Attention [34]. The maximum image resolution was constrained to 384×384 during training to manage memory and allow for larger batch sizes. The whole training takes about 32 hours in this setting.
@@ B.3 Detailed Fine-grained Evaluation Dataset and Metrics
  • Winoground [14]: Tests compositional reasoning abilities through 400 carefully designed instances that require understanding nuanced relationships between objects in images and their textual descriptions.
  • NaturalBench [15]: An expanded benchmark containing 1,200 instances that builds upon Winoground’s principles with greater diversity.
  • MMVP [16]: Comprises 135 instances distributed across 9 distinct visual reasoning categories: Orientation and Direction, Presence of Specific Features, State and Condition, Quantity and Count, Positional and Relational Context, Color and Appearance, Structural and Physical Characteristics, Text, and Viewpoint and Perspective.
  • BiVLC [17]: A large-scale benchmark with 2,933 instances, each categorized into one of three transformation types (Replace, Swap, or Add) that test different aspects of text-image alignment.
@@ Appendix D Related Work
Cross-modal retrieval aims to establish correspondences between visual and textual information in a shared embedding space for effective search [35]. While contrastive learning approaches [1, 36, 37] have advanced the field by training separate encoders on paired datasets, they consistently face challenges from the modality gap [2, 3, 4, 38, 39]. Our work introduces a unified metric that quantifies this gap across different architectures, showing that MLLMs inherently possess strong cross-modal alignment capabilities. Recent MLLM-based retrieval models [8, 10, 9, 21, 40] have reduced this gap through unified architectures, but still rely on relatively simple alignment strategies that limit their effectiveness in tasks requiring nuanced cross-modal understanding.
@@
Preference learning has become a central technique in fine-tuning large language models, particularly in alignment with human feedback [41, 42]. Direct Preference Optimization (DPO) [11] has emerged as a principled alternative to reinforcement learning-based methods, offering a stable, reward-free approach for modeling preferences. While DPO has shown success in natural language domains, its application to cross-modal representation learning remains underexplored. In this work, we extend DPO to retrieval-based settings and propose a reinforcement learning framework that transfers the fine-grained modality alignment priors of MLLMs into the cross-modal representations learned by an MLLM-based retriever.

### v3

@@ 4.3 Ablation Studies
OpenCLIP (ViT-H/14[schuhmann2022laion] 23.64 0.50 47.28 18.05 0.50 36.1
Qwen2-VL-7B [wang2024qwen2] 6.89 6.34 1.09 7.63 2.13 3.58
@@ Appendix A Preliminaries
Recent works [nogueira2020document, sun2023chatgpt] prompt LLMs to make binary relevance judgments and construct candidate rankings. Inspired by these works, we leverage MLLMs to generate fine-grained alignment scores. For each image-text pair, we construct a specific relevance prompt “<image> Does the image align with the text <text>? Answer Yes or No”. This prompt, containing both the image and the text, is fed into the MLLM. The MLLM then processes the input and generates token logits, including those for the target tokens “Yes” and “No”.
@@ B.1 Detailed Training Data Curation
Our primary training data originates from the large-scale OpenImage dataset [kuznetsova2020open]. To curate a high-quality dataset suitable for fine-grained learning, we employed the following procedure:
@@
    Clustering and Deduplication: We first applied clustering and deduplication to the Human-verified OpenImage data to reduce redundancy and group similar images. Following the SemDeDup method [abbas2023semdedup], we set the number of clusters to 50,000 and used an epsilon value of 0.07 to filter out near-duplicate images.
@@
    Hard Negative Mining: For each anchor image in the deduplicated set, we identified hard negative images by retrieving the top-3 most similar images based on their DINOv2 [oquab2023dinov2] embeddings.
@@
    Caption Generation: For each anchor image and its selected top-3 hard negative images, we employed Qwen2.5-VL-72B [bai2025qwen2] to generate corresponding captions for paired images (constructing 6 distinct image pairs from each anchor image and its top-3 hard negative images). To enhance descriptive diversity, we configured the generation temperature to 0.7 and applied the generation process three times, resulting in rich and varied textual representations. As demonstrated in Figure 4, this method enabled us to generate detailed comparative descriptions that effectively capture the visual distinctions between images.
@@ B.2 Detailed Implementation Settings
MAPLE consists of two models: a reward model (Qwen2-VL-7B) and a policy model (initialized with Qwen2-VL-2B or Qwen2-VL-7B). We fine-tune the policy model using Low-Rank Adaptation (LoRA) [hu2022lora], applying it only to the LLM component’s attention and projection layers, while keeping the vision encoder and connector frozen. We employ a differential learning rate strategy, with the base learning rate applied to the LoRA parameters and amplifying the base learning rate by 100 times for the newly added parameters $\tau$ and $\beta$. Key model-related hyperparameters are summarized in Table 5.
@@
Training was conducted on a cluster of 32 NVIDIA A100 GPUs (80GB memory each). To optimize computational efficiency and memory usage, we employed bfloat16 mixed-precision training, gradient checkpointing for the LLM component, and Flash Attention [dao2022flashattention]. The maximum image resolution was constrained to 384×384 during training to manage memory and allow for larger batch sizes. The whole training takes about 32 hours in this setting.
@@ B.3 Detailed Fine-grained Evaluation Dataset and Metrics
  • Winoground [thrush2022winoground]: Tests compositional reasoning abilities through 400 carefully designed instances that require understanding nuanced relationships between objects in images and their textual descriptions.
  • NaturalBench [li2024naturalbench]: An expanded benchmark containing 1,200 instances that builds upon Winoground’s principles with greater diversity.
  • MMVP [tong2024eyes]: Comprises 135 instances distributed across 9 distinct visual reasoning categories: Orientation and Direction, Presence of Specific Features, State and Condition, Quantity and Count, Positional and Relational Context, Color and Appearance, Structural and Physical Characteristics, Text, and Viewpoint and Perspective.
  • BiVLC [miranda2024bivlc]: A large-scale benchmark with 2,933 instances, each categorized into one of three transformation types (Replace, Swap, or Add) that test different aspects of text-image alignment.
@@ Appendix D Related Work
Cross-modal retrieval aims to establish correspondences between visual and textual information in a shared embedding space for effective search [karpathy2015deep]. While contrastive learning approaches [radford2021learning, jia2021scaling, li2021align] have advanced the field by training separate encoders on paired datasets, they consistently face challenges from the modality gap [liang2022mind, jiang2023understanding, schrodi2024two, eslami2024mitigate, ma2024bridging]. Our work introduces a unified metric that quantifies this gap across different architectures, showing that MLLMs inherently possess strong cross-modal alignment capabilities. Recent MLLM-based retrieval models [jiang2024e5, ouali2024discriminative, lin2024mm, lee2024nv, zhou2024megapairs] have reduced this gap through unified architectures, but still rely on relatively simple alignment strategies that limit their effectiveness in tasks requiring nuanced cross-modal understanding.
@@
Preference learning has become a central technique in fine-tuning large language models, particularly in alignment with human feedback [ouyang2022training, bai2022training]. Direct Preference Optimization (DPO) [rafailov2023direct] has emerged as a principled alternative to reinforcement learning-based methods, offering a stable, reward-free approach for modeling preferences. While DPO has shown success in natural language domains, its application to cross-modal representation learning remains underexplored. In this work, we extend DPO to retrieval-based settings and propose a reinforcement learning framework that transfers the fine-grained modality alignment priors of MLLMs into the cross-modal representations learned by an MLLM-based retriever.